OpenStudy (anonymous):

The question is integrate x^3 sqrt(x^2+1)

OpenStudy (blockcolder):

Write the integrand like this: \[\int x \cdot ((x^2+1)-1)\sqrt{x^2+1}\ dx\] Then let u=x^2+1.

OpenStudy (anonymous):

u=x^2+1 du=2xdx x^2=u-1 du/2=xdx new integral: 1/2(u-1)sqrt(u)du =1/2(u-1)(u^1/2)du distribute and integrate normally

OpenStudy (anonymous):

substitute back for u.

OpenStudy (anonymous):

dont we need to eliminate x^3?

OpenStudy (anonymous):

that's the (u-1) = x^2 the du eliminates the other x because du/2=xdx so that accounts for all 3 x's

OpenStudy (anonymous):

(u-1)*du/2=x^3dx

OpenStudy (anonymous):

so whats your answer? because the answer here is \[1/5 (x^{2} + 1)^{3/2} - 1/3 (x^{2} + 1)^{3/2} + C\]

OpenStudy (anonymous):

did u get this answer?

OpenStudy (anonymous):

yes I got this. except double check the answer you're looking at. the first power there should be 5/2 not 3/2 the second one is still 3/2

OpenStudy (anonymous):

yes yessss. im sorry typo. so after how did u integrate the rest?

OpenStudy (anonymous):